OpenStudy (anonymous):

sqrt24-3sqrt45+2sqrt20

OpenStudy (nenadmatematika):

\[2\sqrt{6}-9\sqrt{5}+4\sqrt{5}=2\sqrt{6}-5\sqrt{5}\]